### v1

# Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain

Abstract
In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences such as language understanding. Therefore, building efficient and general-purpose visual backbones based on SSMs is a promising direction. Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive. To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective. We use Fast Fourier Transform (FFT) to obtain the spectrum of the feature map and add it to the original feature map, enabling ViM to model a unified visual representation in both frequency and spatial domains. The introduction of frequency domain information enables ViM to have a global receptive field during scanning. We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains. Moreover, we question the necessity of position embedding in ViM and remove it accordingly in Vim-F, which helps to fully utilize the efficient long-sequence modeling capability of ViM. Finally, we redesign a patch embedding for Vim-F, leveraging a convolutional stem to capture more local correlations, further improving the performance of Vim-F. Code is available at: https://github.com/yws-wxs/Vim-F.
1 Introduction
Visual representation learning stands as a pivotal research topic within the realm of computer vision. In the early era of deep learning, Convolutional Neural Networks (CNNs) occupied a dominant position. CNNs have several built-in inductive biases that make them well-suited to a wide variety of computer vision applications. In addition, CNNs are inherently efficient because when used in a sliding-window manner, the computations are shared. Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks. Except for the initial "patch embedding" module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible. Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism. However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training. The Visual Mamba Models (ViMs) Zhu et al., (2024); Liu et al., (2024); Huang et al., (2024); Pei et al., (2024), which are inspired by recently proposed SSMs Gu and Dao, (2023); Mehta et al., (2023), utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear. Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing. However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships. Vim Zhu et al., (2024) addresses this issue by scanning in bidirectional horizontal directions, while VMamba Liu et al., (2024) adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions. Subsequent works, such as LocalMamba Huang et al., (2024) and EfficientVMamba Pei et al., (2024), have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view. We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field. These efforts have prompted us to focus on the value of scanning in the frequency domain. We know that the use of 2D Discrete Fourier transform (DFT) to convert feature maps into spectrograms does not alter their shapes. The value $P(x,y)$ at any point $(x,y)$ in the spectrogram depends on the entire original feature map. Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field. On the other hand, without considering spectral shifts, the low-frequency components are located in the corners of the spectrogram, whereas the high-frequency components are centered. Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling. Therefore, we adopt the most direct solution, which is to directly add the feature map with its amplitude spectrum, and then let the Vision Mamba encoder learn the fused semantic information. Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance. To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model. Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding three lines of code. Based on Vim-F, we improve the patch embedding to capture more local dependencies. Many successful lightweight transformer models, such as EfficientFormer Li et al., (2022) and MobileViT Mehta and Rastegari, (2022), start their networks with a convolutional stem as patch embedding. This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information. In addition, CNNs are less sensitive to data augmentation, and the training process of a hybrid architecture is more stable. Inspired by these works, we design a patch embedding for Vim-F. Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps. Vim-F that employs this patch embedding module is referred to as Vim-F(H). Finally, we believe that the Mamba model is essentially still a recurrent neural network (RNN), so Vim lacks justification for using position embedding, but removing it will significantly reduce Vim’s performance. Conversely, the use of position embedding has less impact on the performance of Vim-F(H), so our Vim-F(H) does not use position embedding. We outline our contributions below:
  • Based on Vim Zhu et al., (2024), we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships. To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
  • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it. The Vim-F with this patch embedding is referred to as Vim-F(H). Additionally, we question the necessity of position embedding in Vim, thus Vim-F(H) does not utilize position embedding.
  • Through extensive experiments on image classification, object detection, and instance segmentation tasks, the simple Vim variants, Vim-F(H), achieve better performance while maintaining comparable computational costs and parameters. Our work demonstrates that the pure Mamba encoder can benefit from scanning in the frequency domain and introducing more 2D local correlation during the image flattening stage.
2 Related works
2.1 Generic vision backbone design
Nowdays, the realm of vision tasks has been predominantly governed by CNNs and ViTs. Notable CNNs such as VGGNet Simonyan and Zisserman, (2014), Inception Szegedy et al., (2015), ResNet He et al., (2016),ResNeXt Xie et al., (2017), DenseNet Huang et al., (2017), MobileNet Howard et al., (2017), EfficientNet Tan and Le, (2019), and RegNet Radosavovic et al., (2020) have focused on diverse aspects of accuracy, efficiency, and scalability, while promoting numerous influential design principles that have gained widespread acceptance. Around the same time, the network design philosophy for natural language processing took a completely different path, with Transformers replacing Recurrent Neural Networks (RNNs) to become the dominant backbone architecture. Despite the significant differences in the tasks of interest between the language and vision domains, it is noteworthy that the two streams unexpectedly converged. This convergence occurred due to the introduction of ViTs Dosovitskiy et al., (2021), which revolutionized the landscape of network architecture design. The biggest challenge for ViT comes from the design of global attention, which has quadratic complexity relative to the input size. Hierarchical Transformers employ a hybrid approach to solve this problem. For example, the “sliding window” strategy, which involves attention within local windows, has been reintroduced to Transformers Liu et al., (2021), allowing them to behave more similarly to CNNs. Moreover, other studies propose to harness the advantages of CNNs, such as introducing convolution operations, designing hybrid architectures by combining CNN and ViT modules Dai et al., (2021); Srinivas et al., (2021). State Space Models (SSMs) exhibit linear or near-linear scaling complexity with sequence length, making them particularly suitable for handling long sequences. Therefore, SSMs have been widely applied to a range of vision tasks. These applications extend from developing generic foundation models to advancing fields in image segmentation Wang et al., (2024); Ruan and Xiang, (2024), demonstrating the model’s adaptability and potential in visual domain.
2.2 Fourier transforms in neural networks
Because ordinary multiplication in the frequency domain corresponds to a convolution in the time domain, Fast Fourier Transforms (FFTs) have previously been used to approximate or speed up computations in Convolutional Neural Networks Pratt et al., (2017), Recurrent Neural Networks Zhang et al., (2018), and Transformers Tamkin et al., (2020). FNetLee-Thorp et al., (2022) replaces the self-attention sublayer in the Transformer encoder with a standard, non-parametric Fourier transform, achieving an excellent trade-off between speed, memory usage, and accuracy. This suggests that the Fourier transform can serve as an alternative mixing mechanism for hidden representations. GFNet Rao et al., (2021) models the interactions between tokens as a set of learnable global filters that are applied to the spectrum of input features to learn long-range spatial dependencies in the frequency domain with log-linear complexity.
3 Preliminaries
3.1 State Space Models
In mathematics, State Space Models (SSMs) are usually expressed as linear ordinary differential equations (ODEs). These models transform input $D$-dimensional sequence $x(t)\in\mathbb{R}^{L\times D}$ into output sequence $y(t)\in\mathbb{R}^{L\times D}$ by utilizing a learnable latent state $h(t)\in\mathbb{R}^{L\times N}$ that is not directly observable. The mapping process could be denoted as:
$\begin{split}h^{\prime}(t)&=\bm{A}h(t)+\bm{B}x(t),\\ y(t)&=\bm{C}h(t),\end{split}$ (1)
where $\bm{A}\in\mathbb{R}^{N\times N}$, $\bm{B}\in\mathbb{R}^{D\times N}$ and $\bm{C}\in\mathbb{R}^{N\times D}$.
3.2 Discretization
The primary objective of discretization is to convert ODEs into discrete functions, allowing the model to align with the sampling frequency of the input signal for more efficient computation. Following the work Gupta et al., (2022), the continuous parameters ($\bm{A}$, $\bm{B}$) can be discretized by zero-order hold rule with a given sample timescale $\Delta\in\mathbb{R}^{N}$:
$\bar{\bm{A}}=e^{\Delta\bm{A}},$ (2)
$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(e^{\Delta\bm{A}}-\bm{I})\cdot\Delta\bm{B},$
$\bar{\bm{C}}=\bm{C},$
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$
$y(t)=\bar{\bm{C}}h(t),$
where $\bm{\bar{A}}\in\mathbb{R}^{N\times N}$, $\bm{\bar{B}}\in\mathbb{R}^{D\times N}$ and $\bm{\bar{C}}\in\mathbb{R}^{N\times D}$.
To improve computational efficiency, the iterative process described in equation 2 can be accelerated through parallel computing using global convolutional operations:
$\begin{split}\bm{y}&=\bm{x}\circledast\bm{\overline{K}}\\ \text{with}\quad\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{L-1}\bm{\overline{B}}),\end{split}$ (3)
[…]
3.3 Selective State Space Models (S6)
Mamba Gu and Dao, (2023) improves the performance of SSM by introducing Selective State Space Models (S6), allowing the continuous parameters to vary with the input enhances selective information processing across sequences, which extend the discretization process by selection mechanism:
$\bar{\bm{B}}$ $=s_{\bm{B}}(x),$ (4)
$\bar{\bm{C}}$ $=s_{\bm{C}}(x),$
$\Delta$ $=softplus\big{(}\text{Parameter}+s_{\bm{A}}(x)\big{)},$
where $s_{\bm{B}}(x)$ and $s_{\bm{C}}(x)$ are linear functions that project input $x$ into an N-dimensional space, while $s_{\bm{A}}(x)$ broadens a $D$-dimensional linear projection to the necessary dimensions. Parameter represents a trainable parameter matrix. $softplus$ is a mathematical operation defined as: $softplus(x)=log(1+e^{x})$.
Figure 1: The overview of the Vim model.
3.4 Overview of Vim
As shown in Figure 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch. The effective sequence length for Vim is therefore $n=hw/p^{2}$. Then, Vim linearly projects $\mathbf{X}_{p}$ into vectors of size $d$ and adds them to the position embeddings $E_{pos}\in\mathbb{R}^{(n+1)\times d}$, as follows:
$\mathbf{T}_{0}$ $=[\mathbf{X}_{\mathtt{cls}};\mathbf{X}_{p}^{1}\mathbf{W};\mathbf{X}_{p}^{2}\mathbf{W};\cdots;\mathbf{X}_{p}^{i}\mathbf{W};\cdots;\mathbf{X}_{p}^{d}\mathbf{W}]+\mathbf{E}_{pos},$ (5)
where $\mathbf{X}_{p}^{i}$ is the $i$-th patch of $\mathbf{X}$, $\mathbf{W}\in\mathbb{R}^{(p^{2}\cdot c)\times d}$ is the learnable projection matrix. Vim uses class token to represent the whole patch sequence, which is denoted as $\mathbf{X}_{\mathtt{cls}}$. Vim then sends the token sequence ($\mathbf{T}_{l-1}$) to the $l$-th layer of the Vim encoder, and gets the output $\mathbf{T}_{l}$. Finally, Vim normalizes the output class token $\mathbf{T}_{L}^{0}$ and feeds it to the multi-layer perceptron (MLP) head to get the final prediction $\hat{p}$, as follows:
$\mathbf{T}_{l}$ $=\mathbf{Vim}(\mathbf{T}_{l-1})+\mathbf{T}_{l-1},$ (6)
$\mathbf{\it f}$ $=\mathbf{Norm}(\mathbf{T}_{L}^{0}),$
$\hat{p}$ $=\mathbf{MLP}(\mathbf{\it f}),$
where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of layers, and $\mathbf{Norm}$ is the normalization layer.
4 Methodology
In this section, we commence with a thorough introduction to the scanning strategy of Vim-F. Subsequently, we present the specific details of the patch embedding in Vim-F(H). Finally, we analyze the efficiency of Vim-F(H). Through these steps, we aim to provide a comprehensive understanding of Vim-F(H).
4.1 2D scanning in the frequency domain
Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains. Due to the translation invariance of the Fourier transform, the Fourier transforms of figures (a) and (b) are identical.
As shown in Figure 1, the flattening method used in Vim disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens. We address this limitation by introducing frequency-domain scanning. Specifically, we first perform a Fourier transform on the feature map to obtain:
$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}f(x,y)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)},$ (7)
where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively. $f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain. $|F(u,v)|\in\mathbb{R}^{H\times W}$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum. Fourier transform has translation invariance. In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function. This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Figure 2. The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy. Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy. Finally, we use two trainable parameters as the superposition coefficients of the feature map and the corresponding spectrogram to dynamically adjust the information intensity in the spatial and frequency domains. The Mamba encoder establishes a unified representation of visual information in the spatial and frequency domains.
Figure 3: The overall architecture of patch embedding in Vim-F(H).
4.2 Patch embedding
The overall architecture of the patch embedding module is shown in Figure 3.
4.2.1 Convolutional stem
Table 1: Convolutional stem for patch embedding.
Output size
Vim-Ti Vim-S
Block1 56$\times$56 7$\times$7, 48, stride 4 7$\times$7, 48, stride 4
padding 3 padding 3
Block2 28$\times$28 $\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
Block3 14$\times$14 $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
In Vim and ViTs, the patching strategy typically involves using large convolutional kernels (such as kernel sizes of 14 or 16) with non-overlapping convolutions. However, VMamba and Swin Transformer utilize non-overlapping convolutions with a kernel size of 4, indicating that current ViMs haven’t undergone much targeted design in terms of patching. We believe that non-overlapping convolutions have fewer inductive biases, and the self-attention mechanism of the Transformer architecture excels at modeling relationships between patches. Therefore, adopting non-overlapping convolutions can benefit the performance of Transformers. On the contrary, the additional correlation between patches introduced by overlapping convolutions can be advantageous for the scanning mechanism of ViMs.
Based on the above analysis, we can try to explain why Vim uses the position embedding. The bidirectional horizontal scanning strategy employed by Vim makes adjacent tokens in horizontal and vertical directions far apart in the sequence. In addition, the lack of data correlation between tokens makes it difficult for the model to understand the spatial relationship in the vertical direction. The introduction of the trainable position embedding can explicitly provide the model with the absolute position of each token in the plane space, thus alleviating the above problems. In contrast, it is simpler and more flexible to remove the position embedding and establish spatial correlation through overlapping convolutions, enabling the model to learn spatial positional relationships.
Specifically, we first utilize a $7\times 7$ convolutional layer with a stride of 4 to model the local correlation of the input image. Then, we downsample the feature map twice through two downsampling blocks. Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer. The details of the adaptation with the two base models Zhu et al., (2024) (Vim-Ti and Vim-S) are shown in Table 1 and Figure 3.
4.3 Efficiency analysis
Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation. The Cooley-Tukey FFT algorithm decomposes a DFT of a sequence with length N into two DFTs of sub-sequences with length N/2, significantly reducing the computational complexity. During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results. For a 2D-DFT with dimensions $(H,W)$, its complexity is approximately $O(H\cdot W\cdot log(H)\cdot log(W))$. Compared to the overall computational complexity of the model, the computational cost of this part can be negligible.
The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16. The convolutional stem for patch embedding in Vim-F(H) comprises more convolutional layers, but the computational cost increases moderately due to the use of smaller convolutional kernels. Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035. In summary, there is no significant difference in computational cost between Vim-F(H) and Vim. LocalVimHuang et al., (2024) is one of the representative works that improve the scanning strategy for Vim. In terms of efficiency, LocalVim’s selective scanning branches are twice as many as our Vim-F(H), and it uses a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in the features of each branch. Therefore, our Vim-F(H) achieves a better balance between performance and efficiency.
Table 2: Comparison of different backbones on ImageNet-1K classification. :$\dagger$:† : We are unable to reproduce the results through the source code provided by the author, and one possible reason is the significant difference in hardware and the use of linear scaling rule. The results in () are from the original paper.
Method Params (M) FLOPs (G) Top-1 (%)
PVTv2-B0 Wang et al., (2022) 3 0.6 70.5
DeiT-Ti Touvron et al., (2021) 6 1.3 74.4
MobileViT-XS Mehta and Rastegari, (2021) 2 1.0 74.8
LVT Yang et al., (2022) 6 0.9 74.8
[…]
Vim-Ti-F(H) (ours) 7 1.5 76.0
LocalVim-T Huang et al., (2024) 8 1.5 76.5
RegNetY-1.6G Radosavovic et al., (2020) 11 1.6 78.0
MobileViT-S Mehta and Rastegari, (2021) 6 2.0 78.4
EfficientVMamba-S Pei et al., (2024) 11 1.3 78.7
DeiT-S Touvron et al., (2021) 22 4.6 79.8
RegNetY-4G Radosavovic et al., (2020) 21 4.0 80.0
Swin-T Liu et al., (2021) 29 4.5 81.3
[…]
Vim-S-F(H) (ours) 26 5.1 80.4
LocalVim-SHuang et al., (2024) 28 4.8 81.0
EfficientVMamba-BPei et al., (2024) 33 4.0 81.8
VMamba-T Liu et al., (2024) 22 5.6 82.2
5 Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation.
5.1 Image classification
Settings. We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224\times 224$ input images, we employ AdamW with a momentum of 0.9 and a weight decay of 0.05 to optimize models. During testing, we apply a center crop on the validation set to crop out $224\times 224$ images. We train the Vim-F(H) models for 300 epochs using a cosine schedule and EMA. Unlike Vim, our experiments are performed on 3 A6000 GPUs. Therefore, we adjusted the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling ruleGoyal et al., (2017).
Results. We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table 2. Compared to our seminal Vim, our method has made substantial progress, with Vim-Ti-F(H) and Vim-S-F(H) achieving 1.3% and 0.8% higher accuracy than Vim-Ti and Vim-S, respectively. It is worth noting that SOTA ViMs such as VMamba Liu et al., (2024), LocalMamba Huang et al., (2024), and EfficientVMamba Pei et al., (2024) all adopt more complex hybrid encoding architectures, even introducing attention mechanisms. However, Vim-F still uses a pure SSM encoder to model visual information. Vim-Ti-F(H) achieves superior accuracy compared to highly optimized lightweight ViT variants, including DeiT-TiTouvron et al., (2021), MobileViT-XS Mehta and Rastegari, (2021), and LVT Yang et al., (2022), while also narrowing the gap in performance between Vim and other leading ViMs.
Long sequence fine-tuning. ViM with a pure Mamba encoder can compress signals at a low cost and reconstruct them into longer sequences. For visual tasks, flattening images into longer 1D sequences and fine-tuning the model can allow the model to capture more details and improve performance. Due to the use of absolute position embedding, Vim needs to expand the encoding length through interpolation methods. In addition, because the position embedding is located at the front of the entire network, more epochs of fine-tuning are required for adaptation. We set the stride of the first convolutional layer in the patch embedding to 2, and set the learning rate to $1\times 10^{-5}$ for 1 epoch of fine-tuning. Vim-Ti-F(H) can achieve a 0.8% increase in accuracy.
Table 3: The classification result of Vim-F(H) on ImageNet-1K after 1 epoch of fine-tuning. Different from the pre-training stage, the input image is flattened into 784 tokens (28$\times$28).
Method Params (M) Top-1 (%)
Vim-Ti-F(H) (Fine-tuning for 1 epoch) 7 76.8
Vim-S-F(H) (Fine-tuning for 1 epoch) 26 81.2
Table 4: Object detection and instance segmentation results on COCO.
Mask R-CNN 1$\times$ schedule
[…]
LightViT-T 28M 187G 37.8 60.7 40.4 35.9 57.8 38.0
ResNet-18 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
EfficientVMamba-S 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
ResNet-50 44M 260G 38.2 58.8 41.4 34.7 55.7 37.2
LocalVMamba-T 45M 291G 46.7 68.7 50.8 42.2 65.7 45.5
Swin-T 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
ConvNeXt-T 48M 262G 44.2 66.6 48.3 40.1 63.3 42.8
ViT-Adapter-S 48M 403G 44.7 65.8 48.3 39.9 62.5 42.8
ResNet-101 63M 336G 38.2 58.8 41.4 34.7 55.7 37.2
Swin-S 69M 354G 44.8 66.6 48.9 40.9 63.2 44.2
ConvNeXt-S 70M 348G 45.4 67.9 50.0 41.8 65.2 45.1
Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
Vim-Ti-F(H) (ours) 27M 189G 37.7 60.8 40.9 35.7 57.8 37.9
Vim-S 44M 272G 42.9 65.4 47.1 39.2 62.1 42.1
Vim-S-F(H) (ours) 44M 272G 43.1 65.2 47.3 39.3 62.2 42.3
5.2 Object detection and instance segmentation
Settings. We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F(H) for object detection and instance segmentation on the MSCOCO 2017 dataset. Since the backbone of Vim-F(H) is non-hierarchical, following ViTDetLi et al., (2021), we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector. The remaining settings were consistent with SwinLiu et al., (2021). Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule). The learning rate is initialized at $1\times 10^{-4}$ and is reduced by a factor of 10×10$\times$10 × at the 9th and 11th epochs.
Results. We summarize the comparison results of Vim-F(H) with other backbones in Table 4. It can be seen that our Vim-F(H) consistently outperforms Vim. Similar to the results on classification tasks, Vim-F(H) with a pure SSM encoder achieves a good balance between the number of parameters and computational cost, achieving comparable results with advanced CNNs and ViTs.
6 Conclusion and limitation
One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data. Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships. We observed that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field. We adopted a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains. Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps. We also introduce additional local correlations during patch embedding by using overlapping convolutions. Our Vim-F(H) is motivated by a simple idea and is easy to implement. Extensive experiments have demonstrated that it effectively improves the performance of Vim, achieving competitive results compared to advanced ViTs, CNNs, and ViMs. The effectiveness of our method for ViMs with hybrid encoders has not been fully studied.
Appendix A Appendix
A.1 Ablation study
Effect of frequency domain scanning. We evaluated the scanning strategy of Vim-F, and the experimental details are shown in Table 5. As mentioned earlier, position embeddings are crucial for Vim, and removing them leads to a 1.7% performance drop for Vim-Ti. However, the global receptive field brought by frequency domain scanning reduces the model’s dependence on positional embeddings, resulting in a smaller performance drop for Vim-Ti-F. When both models use positional embeddings, Vim-Ti-F outperforms Vim-Ti by 0.4%.
Effect of convolutional stem. As shown in Table 5, further replacing the traditional patch embedding in Vim-Ti with our proposed convolutional stem leads to a 0.9% improvement in performance. This shows that introducing additional local correlation into the patch embedding is effective.
Table 5: Ablation study of frequency domain scanning and patch embedding in Vim-Ti-F(H) on ImageNet-1K.
Model Top-1 (%)
Vim-Ti (without Position Embedding) 73.0%
Vim-Ti (base) 74.7%
Vim-Ti-F 74.2%
Vim-Ti-F + Position Embedding 75.1%
Vim-Ti-F(H) 76.0%

### v2

@@
 Abstract
-In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences such as language understanding.
-Therefore, building efficient and general-purpose visual backbones based on SSMs is a promising direction.
+In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences.
 Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive.
 To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective.
-We use Fast Fourier Transform (FFT) to obtain the spectrum of the feature map and add it to the original feature map, enabling ViM to model a unified visual representation in both frequency and spatial domains.
-The introduction of frequency domain information enables ViM to have a global receptive field during scanning.
+We believe that the introduction of frequency domain information can enable ViM to achieve a better global receptive field during the scanning process.
 We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains.
@@
 Code is available at: https://github.com/yws-wxs/Vim-F.
-1 Introduction
+Index Terms:
+Vision Mamba, State Space Model, Fast Fourier Transform, Scanning strategy
+I Introduction
 Visual representation learning stands as a pivotal research topic within the realm of computer vision.
@@
 Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks.
-Except for the initial "patch embedding" module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible.
+Except for the initial ”patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible.
 Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism.
 However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
+Figure 1: The overview of the Vim model.
 Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training.
-The Visual Mamba Models (ViMs) Zhu et al., (2024); Liu et al., (2024); Huang et al., (2024); Pei et al., (2024), which are inspired by recently proposed SSMs Gu and Dao, (2023); Mehta et al., (2023), utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear.
+The Visual Mamba Models (ViMs) [1, 2, 3, 4], which are inspired by recently proposed SSMs [5, 6], utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear.
 Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing.
 However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships.
-Vim Zhu et al., (2024) addresses this issue by scanning in bidirectional horizontal directions, while VMamba Liu et al., (2024) adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions.
-Subsequent works, such as LocalMamba Huang et al., (2024) and EfficientVMamba Pei et al., (2024), have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view.
+Vim [1] addresses this issue by scanning in bidirectional horizontal directions, while VMamba [2] adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions.
+Subsequent works, such as LocalMamba [3] and EfficientVMamba [4], have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view.
 We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field.
@@
 Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field.
-On the other hand, without considering spectral shifts, the low-frequency components are located in the corners of the spectrogram, whereas the high-frequency components are centered.
+On the other hand, without considering spectral shifts, the high-frequency components are located in the corners of the spectrogram, whereas the low-frequency components are centered.
 Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling.
-Therefore, we adopt the most direct solution, which is to directly add the feature map with its amplitude spectrum, and then let the Vision Mamba encoder learn the fused semantic information.
+Therefore, we adopt the most direct solution, which is to add the amplitude spectrum of the feature map to the original feature map, allowing the visual Mamba encoder to learn the fused semantic information.
 Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance.
 To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model.
-Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding three lines of code.
-Based on Vim-F, we improve the patch embedding to capture more local dependencies.
-Many successful lightweight transformer models, such as EfficientFormer Li et al., (2022) and MobileViT Mehta and Rastegari, (2022), start their networks with a convolutional stem as patch embedding.
+Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding a few lines of code.
+Secondly, we improve the patch embedding to capture more local dependencies.
+Many successful lightweight transformer models, such as EfficientFormer [7] and MobileViT [8], start their networks with a convolutional stem as patch embedding.
 This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information.
@@
 Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps.
-Vim-F that employs this patch embedding module is referred to as Vim-F(H).
-Finally, we believe that the Mamba model is essentially still a recurrent neural network (RNN), so Vim lacks justification for using position embedding, but removing it will significantly reduce Vim’s performance.
-Conversely, the use of position embedding has less impact on the performance of Vim-F(H), so our Vim-F(H) does not use position embedding.
+Finally, we argue that the Mamba model is essentially still a recurrent neural network (RNN).
+Therefore, Vim lacks the rationale for using positional embeddings, and thus our Vim-F does not employ positional embeddings.
+Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining overall linear complexity.
+We refer to this hybrid architecture as Vim-F(H).
 We outline our contributions below:
-  • Based on Vim Zhu et al., (2024), we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships.
+  • Based on Vim [1], we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships.
 To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
   • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it.
-The Vim-F with this patch embedding is referred to as Vim-F(H).
-Additionally, we question the necessity of position embedding in Vim, thus Vim-F(H) does not utilize position embedding.
-  • Through extensive experiments on image classification, object detection, and instance segmentation tasks, the simple Vim variants, Vim-F(H), achieve better performance while maintaining comparable computational costs and parameters.
-Our work demonstrates that the pure Mamba encoder can benefit from scanning in the frequency domain and introducing more 2D local correlation during the image flattening stage.
-2 Related works
-2.1 Generic vision backbone design
-Nowdays, the realm of vision tasks has been predominantly governed by CNNs and ViTs.
-Notable CNNs such as VGGNet Simonyan and Zisserman, (2014), Inception Szegedy et al., (2015), ResNet He et al., (2016),ResNeXt Xie et al., (2017), DenseNet Huang et al., (2017), MobileNet Howard et al., (2017), EfficientNet Tan and Le, (2019), and RegNet Radosavovic et al., (2020) have focused on diverse aspects of accuracy, efficiency, and scalability, while promoting numerous influential design principles that have gained widespread acceptance.
-Around the same time, the network design philosophy for natural language processing took a completely different path, with Transformers replacing Recurrent Neural Networks (RNNs) to become the dominant backbone architecture.
-Despite the significant differences in the tasks of interest between the language and vision domains, it is noteworthy that the two streams unexpectedly converged.
-This convergence occurred due to the introduction of ViTs Dosovitskiy et al., (2021), which revolutionized the landscape of network architecture design.
-The biggest challenge for ViT comes from the design of global attention, which has quadratic complexity relative to the input size.
-Hierarchical Transformers employ a hybrid approach to solve this problem.
-For example, the “sliding window” strategy, which involves attention within local windows, has been reintroduced to Transformers Liu et al., (2021), allowing them to behave more similarly to CNNs.
-Moreover, other studies propose to harness the advantages of CNNs, such as introducing convolution operations, designing hybrid architectures by combining CNN and ViT modules Dai et al., (2021); Srinivas et al., (2021).
-State Space Models (SSMs) exhibit linear or near-linear scaling complexity with sequence length, making them particularly suitable for handling long sequences.
-Therefore, SSMs have been widely applied to a range of vision tasks.
-These applications extend from developing generic foundation models to advancing fields in image segmentation Wang et al., (2024); Ruan and Xiang, (2024), demonstrating the model’s adaptability and potential in visual domain.
-2.2 Fourier transforms in neural networks
-Because ordinary multiplication in the frequency domain corresponds to a convolution in the time domain, Fast Fourier Transforms (FFTs) have previously been used to approximate or speed up computations in Convolutional Neural Networks Pratt et al., (2017), Recurrent Neural Networks Zhang et al., (2018), and Transformers Tamkin et al., (2020).
-FNetLee-Thorp et al., (2022) replaces the self-attention sublayer in the Transformer encoder with a standard, non-parametric Fourier transform, achieving an excellent trade-off between speed, memory usage, and accuracy.
-This suggests that the Fourier transform can serve as an alternative mixing mechanism for hidden representations.
-GFNet Rao et al., (2021) models the interactions between tokens as a set of learnable global filters that are applied to the spectrum of input features to learn long-range spatial dependencies in the frequency domain with log-linear complexity.
-3 Preliminaries
-3.1 State Space Models
+Additionally, we question the necessity of position embedding in Vim, thus Vim-F does not utilize position embedding.
+  • The experimental results demonstrate that the simple Vim variant, Vim-F, achieves better performance while maintaining similar computational costs and parameters.
+Additionally, introducing linear attention to reduce Vim-F’s sensitivity to the order of input sequences can further enhance the model’s performance.
+II Related Works
+II-A State Space Models
 In mathematics, State Space Models (SSMs) are usually expressed as linear ordinary differential equations (ODEs).
@@
 where $\bm{A}\in\mathbb{R}^{N\times N}$, $\bm{B}\in\mathbb{R}^{D\times N}$ and $\bm{C}\in\mathbb{R}^{N\times D}$.
-3.2 Discretization
-The primary objective of discretization is to convert ODEs into discrete functions, allowing the model to align with the sampling frequency of the input signal for more efficient computation.
-Following the work Gupta et al., (2022), the continuous parameters ($\bm{A}$, $\bm{B}$) can be discretized by zero-order hold rule with a given sample timescale $\Delta\in\mathbb{R}^{N}$:
-$\bar{\bm{A}}=e^{\Delta\bm{A}},$ (2)
-$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(e^{\Delta\bm{A}}-\bm{I})\cdot\Delta\bm{B},$
-$\bar{\bm{C}}=\bm{C},$
-$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$
-$y(t)=\bar{\bm{C}}h(t),$
-where $\bm{\bar{A}}\in\mathbb{R}^{N\times N}$, $\bm{\bar{B}}\in\mathbb{R}^{D\times N}$ and $\bm{\bar{C}}\in\mathbb{R}^{N\times D}$.
-To improve computational efficiency, the iterative process described in equation 2 can be accelerated through parallel computing using global convolutional operations:
-$\begin{split}\bm{y}&=\bm{x}\circledast\bm{\overline{K}}\\ \text{with}\quad\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{L-1}\bm{\overline{B}}),\end{split}$ (3)
-[…]
-3.3 Selective State Space Models (S6)
-Mamba Gu and Dao, (2023) improves the performance of SSM by introducing Selective State Space Models (S6), allowing the continuous parameters to vary with the input enhances selective information processing across sequences, which extend the discretization process by selection mechanism:
-$\bar{\bm{B}}$ $=s_{\bm{B}}(x),$ (4)
-$\bar{\bm{C}}$ $=s_{\bm{C}}(x),$
-$\Delta$ $=softplus\big{(}\text{Parameter}+s_{\bm{A}}(x)\big{)},$
-where $s_{\bm{B}}(x)$ and $s_{\bm{C}}(x)$ are linear functions that project input $x$ into an N-dimensional space, while $s_{\bm{A}}(x)$ broadens a $D$-dimensional linear projection to the necessary dimensions.
-Parameter represents a trainable parameter matrix.
-$softplus$ is a mathematical operation defined as: $softplus(x)=log(1+e^{x})$.
-Figure 1: The overview of the Vim model.
-3.4 Overview of Vim
-As shown in Figure 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch.
+II-B Overview of Vim
+As mentioned earlier, the effectiveness of our work is demonstrated through the improvement of Vim.
+Therefore, we provide a brief introduction to the overall architecture of Vim.
+As shown in Fig.
+1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch.
 The effective sequence length for Vim is therefore $n=hw/p^{2}$.
@@
 $\hat{p}$ $=\mathbf{MLP}(\mathbf{\it f}),$
-where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of layers, and $\mathbf{Norm}$ is the normalization layer.
-4 Methodology
-In this section, we commence with a thorough introduction to the scanning strategy of Vim-F.
-Subsequently, we present the specific details of the patch embedding in Vim-F(H).
-Finally, we analyze the efficiency of Vim-F(H).
-Through these steps, we aim to provide a comprehensive understanding of Vim-F(H).
-4.1 2D scanning in the frequency domain
+where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of blocks, and $\mathbf{Norm}$ is the normalization layer.
 Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains.
 Due to the translation invariance of the Fourier transform, the Fourier transforms of figures (a) and (b) are identical.
-As shown in Figure 1, the flattening method used in Vim disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens.
+III Rethinking the Scanning Strategy of ViMs
+The non-causal nature of two-dimensional spatial patterns in images is inherently inconsistent with the causal processing framework of SSM.
+Therefore, it is meaningful to explore methods for serializing images.
+However, the majority of existing work on scanning strategies is typically based on a rather idealized assumption that introducing more or more complex inductive biases regarding the order of one-dimensional sequences can potentially enhance Mamba’s ability to understand local image correlations.
+Theoretically, regardless of the order, reducing the distance between adjacent tokens in one direction in two-dimensional space inevitably increases the distance between adjacent tokens in another direction, thus these strategies are essentially equivalent.
+Additionally, in the absence of global information serving as a ”frame of reference,” these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences.
+In practice, works involving scanning strategy research, including Vim [1], VMamba [2], LocalMamba [3], EfficientVMamba [4], and PlainMamba [9], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies.
+Zhu et al.
+[10] conduct a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images.
+Through extensive experiments on the LoveDA, ISPRS Potsdam, and ISPRS Vaihingen datasets, they demonstrate that no single scanning strategy outperforms others, regardless of their complexity or the number of scanning directions involved.
+In summary, an effective scanning strategy needs to maintain a global receptive field throughout the scanning process and demonstrate its effectiveness through analysis and comparative experiments.
+IV Methodology
+In this section, we first elaborate on the significance and benefits of frequency-domain scanning.
+Subsequently, we present the specific details of patch embedding in Vim-F.
+Finally, we discuss the composition of the Vim-F and the Vim-F(H) block.
+IV-A 2D Scanning in the Frequency Domain
+The flattening method used in Vim [1] disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens.
 We address this limitation by introducing frequency-domain scanning.
@@
 In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function.
-This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Figure 2.
+This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Fig.
+2.
 The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy.
 Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy.
-Finally, we use two trainable parameters as the superposition coefficients of the feature map and the corresponding spectrogram to dynamically adjust the information intensity in the spatial and frequency domains.
-The Mamba encoder establishes a unified representation of visual information in the spatial and frequency domains.
-Figure 3: The overall architecture of patch embedding in Vim-F(H).
-4.2 Patch embedding
-The overall architecture of the patch embedding module is shown in Figure 3.
-4.2.1 Convolutional stem
-Table 1: Convolutional stem for patch embedding.
-Output size
-Vim-Ti Vim-S
-Block1 56$\times$56 7$\times$7, 48, stride 4 7$\times$7, 48, stride 4
-padding 3 padding 3
-Block2 28$\times$28 $\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
-Block3 14$\times$14 $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
+IV-B Patch Embedding
 In Vim and ViTs, the patching strategy typically involves using large convolutional kernels (such as kernel sizes of 14 or 16) with non-overlapping convolutions.
@@
 Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer.
-The details of the adaptation with the two base models Zhu et al., (2024) (Vim-Ti and Vim-S) are shown in Table 1 and Figure 3.
-4.3 Efficiency analysis
+The details of the adaptation with the two base models [1] (Vim-Ti and Vim-S) are shown in Table I.
+TABLE I: Convolutional stem for patch embedding.
+Output size
+Vim-Ti Vim-S
+Block1 56$\times$56 7$\times$7, 48, stride 4 7$\times$7, 48, stride 4
+padding 3 padding 3
+Block2 28$\times$28 $\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
+Block3 14$\times$14 $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
+The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16.
+The convolutional stem for patch embedding in Vim-F comprises more convolutional layers, but the additional computational cost is negligible due to the use of smaller convolutional kernels.
+Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035.
+IV-C Vim-F Block
+As mentioned earlier, the Vim-F block is an improvement over the Vim block, as shown in Fig.
+3.
+Specifically, a 2D FFT is performed on the input to obtain the spectrogram.
+Then, the spectrogram is added to the original input, with the intensity of the frequency and spatial domain information being adjusted by two trainable parameters, $\alpha$ and $\beta$, respectively.
+After the linear layer mixes the hidden dimensions of the feature map, the cross-domain information is sent to the Vim block for processing.
+Figure 3: Vim-F block
 Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation.
@@
 During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results.
-For a 2D-DFT with dimensions $(H,W)$, its complexity is approximately $O(H\cdot W\cdot log(H)\cdot log(W))$.
-Compared to the overall computational complexity of the model, the computational cost of this part can be negligible.
-The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16.
-The convolutional stem for patch embedding in Vim-F(H) comprises more convolutional layers, but the computational cost increases moderately due to the use of smaller convolutional kernels.
-Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035.
-In summary, there is no significant difference in computational cost between Vim-F(H) and Vim.
-LocalVimHuang et al., (2024) is one of the representative works that improve the scanning strategy for Vim.
-In terms of efficiency, LocalVim’s selective scanning branches are twice as many as our Vim-F(H), and it uses a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in the features of each branch.
-Therefore, our Vim-F(H) achieves a better balance between performance and efficiency.
-Table 2: Comparison of different backbones on ImageNet-1K classification.
-:$\dagger$:† : We are unable to reproduce the results through the source code provided by the author, and one possible reason is the significant difference in hardware and the use of linear scaling rule.
-The results in () are from the original paper.
+For a 2D-DFT with dimensions $(L,D)$, its complexity is $\Omega(L\cdot D\cdot logL\cdot logD)$.
+The feedforward network is implemented by a linear layer, with a complexity of $\Omega(L\cdot D^{2})$.
+Therefore, compared to the Vim block, the additional overhead of the Vim-F block is approximately linearly related to the sequence length $L$.
+IV-D Vim-F(H) Block
+Observing Eq.
+(1), it is evident that since all elements of matrix $A$ are strictly confined within the range of 0 to 1 [5], the early information stored in the hidden state $h$ will continuously decay.
+This leads to Mamba being sensitive to the order of the input sequence.
+We propose that by weighting the input sequence with attention mechanisms before processing with Mamba, this strong local bias can be mitigated, thereby enhancing Mamba’s capability to handle long sequences.
+We integrate the Vim-F block with the linear attention [11] to maintain approximately linear complexity, which we term Vim-F(H), and its structure is depicted in Fig.
+4.
+Figure 4: Vim-F(H) block
+TABLE II: Comparison of different backbones on ImageNet-1K classification.
 Method Params (M) FLOPs (G) Top-1 (%)
-PVTv2-B0 Wang et al., (2022) 3 0.6 70.5
-DeiT-Ti Touvron et al., (2021) 6 1.3 74.4
-MobileViT-XS Mehta and Rastegari, (2021) 2 1.0 74.8
-LVT Yang et al., (2022) 6 0.9 74.8
-[…]
-Vim-Ti-F(H) (ours) 7 1.5 76.0
-LocalVim-T Huang et al., (2024) 8 1.5 76.5
-RegNetY-1.6G Radosavovic et al., (2020) 11 1.6 78.0
-MobileViT-S Mehta and Rastegari, (2021) 6 2.0 78.4
-EfficientVMamba-S Pei et al., (2024) 11 1.3 78.7
-DeiT-S Touvron et al., (2021) 22 4.6 79.8
-RegNetY-4G Radosavovic et al., (2020) 21 4.0 80.0
-Swin-T Liu et al., (2021) 29 4.5 81.3
-[…]
-Vim-S-F(H) (ours) 26 5.1 80.4
-LocalVim-SHuang et al., (2024) 28 4.8 81.0
-EfficientVMamba-BPei et al., (2024) 33 4.0 81.8
-VMamba-T Liu et al., (2024) 22 5.6 82.2
-5 Experiments
+PVTv2-B0 [12] 3 0.6 70.5
+DeiT-Ti [13] 6 1.3 74.4
+MobileViT-XS [14] 2 1.0 74.8
+LVT [15] 6 0.9 74.8
+Vim-Ti [1] 7 1.5 76.1
+LocalVim-T [3] 8 1.5 76.5
+Vim-Ti-F (ours) 8 1.8 76.7
+PlainMamba-L1 [9] 7 3.0 77.9
+RegNetY-1.6G [16] 11 1.6 78.0
+Vim-Ti-F(H) (ours) 9 2.1 78.2
+MobileViT-S [14] 6 2.0 78.4
+EfficientVMamba-S [4] 11 1.3 78.7
+DeiT-S [13] 22 4.6 79.8
+RegNetY-4G [16] 21 4.0 80.0
+Vim-S [1] 26 5.1 80.5
+Vim-S-F (ours) 28 6.1 80.9
+LocalVim-S[3] 28 4.8 81.0
+Swin-T [17] 29 4.5 81.3
+Vim-S-F(H) (ours) 37 6.8 81.6
+EfficientVMamba-B[4] 33 4.0 81.8
+VMamba-T [2] 22 5.6 82.2
+V Experiments
 This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation.
-5.1 Image classification
+The specific architectures of Vim-F and Vim-F(H) are as follows: We replaced only the first 6 Vim blocks (25% of the total number of blocks) with either Vim-F or Vim-F(H) blocks to balance performance and efficiency while maintaining approximately the same number of parameters.
+The number of linear attention heads in Vim-F(H) is set to 24.
+V-A Image Classification
 Settings.
@@
 For fair comparisons, our training settings mainly follow Vim.
-Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations.
-When training on $224\times 224$ input images, we employ AdamW with a momentum of 0.9 and a weight decay of 0.05 to optimize models.
-During testing, we apply a center crop on the validation set to crop out $224\times 224$ images.
-We train the Vim-F(H) models for 300 epochs using a cosine schedule and EMA.
-Unlike Vim, our experiments are performed on 3 A6000 GPUs.
-Therefore, we adjusted the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling ruleGoyal et al., (2017).
+Due to hardware limitations, our experiments are performed on 3 A6000 GPUs.
+Therefore, we adjust the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling rule[18].
 Results.
-We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table 2.
-Compared to our seminal Vim, our method has made substantial progress, with Vim-Ti-F(H) and Vim-S-F(H) achieving 1.3% and 0.8% higher accuracy than Vim-Ti and Vim-S, respectively.
-It is worth noting that SOTA ViMs such as VMamba Liu et al., (2024), LocalMamba Huang et al., (2024), and EfficientVMamba Pei et al., (2024) all adopt more complex hybrid encoding architectures, even introducing attention mechanisms.
+We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table II.
+Vim-F outperforms the original Vim in terms of performance, and Vim-F(H) further enhances Vim-F’s performance by introducing linear attention.
+It is worth noting that SOTA ViMs such as VMamba [2], PlainMamba [9], and EfficientVMamba [4] all adopt more complex CNN-Mamba hybrid encoding architectures, and their network structures are similar to advanced convolutional networks like ConvNeXt [19].
+Since these works have not conducted the necessary comparative experiments to address this issue, it is reasonable to suspect that the excellent performance of these ViMs may primarily stem from the use of depthwise separable convolutions.
 However, Vim-F still uses a pure SSM encoder to model visual information.
-Vim-Ti-F(H) achieves superior accuracy compared to highly optimized lightweight ViT variants, including DeiT-TiTouvron et al., (2021), MobileViT-XS Mehta and Rastegari, (2021), and LVT Yang et al., (2022), while also narrowing the gap in performance between Vim and other leading ViMs.
-Long sequence fine-tuning.
-ViM with a pure Mamba encoder can compress signals at a low cost and reconstruct them into longer sequences.
-For visual tasks, flattening images into longer 1D sequences and fine-tuning the model can allow the model to capture more details and improve performance.
-Due to the use of absolute position embedding, Vim needs to expand the encoding length through interpolation methods.
-In addition, because the position embedding is located at the front of the entire network, more epochs of fine-tuning are required for adaptation.
-We set the stride of the first convolutional layer in the patch embedding to 2, and set the learning rate to $1\times 10^{-5}$ for 1 epoch of fine-tuning.
-Vim-Ti-F(H) can achieve a 0.8% increase in accuracy.
-Table 3: The classification result of Vim-F(H) on ImageNet-1K after 1 epoch of fine-tuning. Different from the pre-training stage, the input image is flattened into 784 tokens (28$\times$28).
-Method Params (M) Top-1 (%)
-Vim-Ti-F(H) (Fine-tuning for 1 epoch) 7 76.8
-Vim-S-F(H) (Fine-tuning for 1 epoch) 26 81.2
-Table 4: Object detection and instance segmentation results on COCO.
-Mask R-CNN 1$\times$ schedule
+LocalVim[3] is one of the representative works that improve the scanning strategy for Vim.
+LocalVim has designed 8 scanning strategies and employs neural architecture search to determine the scanning strategy used by each of the 4 scanning branches in every LocalVim block.
+Additionally, it utilizes a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in each branch feature.
+Moreover, the total number of blocks in LocalVim has not followed Vim’s 24 but has been reduced to 20.
+In comparison, our works achieve a better balance between performance, efficiency, and model complexity.
+TABLE III: Object detection and instance segmentation results on COCO.
 […]
-LightViT-T 28M 187G 37.8 60.7 40.4 35.9 57.8 38.0
 ResNet-18 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
+Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
+PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
+Vim-F-Ti (ours) 28M 189G 37.7 60.8 40.9 35.7 57.8 38.0
+ResNet-50 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
+Vim-F(H)-Ti (ours) 28M 189G 38.2 61.8 41.7 35.9 58.6 38.0
 EfficientVMamba-S 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
-PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
-ResNet-50 44M 260G 38.2 58.8 41.4 34.7 55.7 37.2
-LocalVMamba-T 45M 291G 46.7 68.7 50.8 42.2 65.7 45.5
+ResNet-101 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
+Vim-S 44M 272G 40.9 63.9 45.1 37.9 60.8 40.7
+Vim-F-S (ours) 47M 273G 41.6 64.1 45.5 38.4 60.9 41.3
 Swin-T 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
-ConvNeXt-T 48M 262G 44.2 66.6 48.3 40.1 63.3 42.8
-ViT-Adapter-S 48M 403G 44.7 65.8 48.3 39.9 62.5 42.8
-ResNet-101 63M 336G 38.2 58.8 41.4 34.7 55.7 37.2
+Vim-F(H)-S (ours) 57M 275G 42.9 65.4 47.1 39.2 62.1 42.1
 Swin-S 69M 354G 44.8 66.6 48.9 40.9 63.2 44.2
-ConvNeXt-S 70M 348G 45.4 67.9 50.0 41.8 65.2 45.1
-Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
-Vim-Ti-F(H) (ours) 27M 189G 37.7 60.8 40.9 35.7 57.8 37.9
-Vim-S 44M 272G 42.9 65.4 47.1 39.2 62.1 42.1
-Vim-S-F(H) (ours) 44M 272G 43.1 65.2 47.3 39.3 62.2 42.3
-5.2 Object detection and instance segmentation
+V-B Object Detection and Instance Segmentation
 Settings.
-We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F(H) for object detection and instance segmentation on the MSCOCO 2017 dataset.
-Since the backbone of Vim-F(H) is non-hierarchical, following ViTDetLi et al., (2021), we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector.
-The remaining settings were consistent with SwinLiu et al., (2021).
+We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F for object detection and instance segmentation on the MSCOCO 2017 dataset.
+Since the backbone of Vim-F is non-hierarchical, following ViTDet[20], we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector.
+The remaining settings were consistent with Swin[17].
 Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule).
@@
 Results.
-We summarize the comparison results of Vim-F(H) with other backbones in Table 4.
-It can be seen that our Vim-F(H) consistently outperforms Vim.
-Similar to the results on classification tasks, Vim-F(H) with a pure SSM encoder achieves a good balance between the number of parameters and computational cost, achieving comparable results with advanced CNNs and ViTs.
-6 Conclusion and limitation
+We summarize the comparison results of Vim-F and Vim-F(H) with other backbones in Table III.
+It can be seen that our Vim-F consistently outperforms Vim.
+Similar to the results in classification tasks, Vim-F(H) exhibits better performance than Vim-F and is competitive with advanced ViT architectures.
+VI Conclusion
 One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data.
 Consequently, scanning one-dimensional tokens with specific rules can limit the model’s ability to accurately interpret spatial relationships.
-We observed that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field.
-We adopted a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains.
+We observe that scanning in the frequency domain, leveraging the translation invariance and global nature of Fourier transforms, can reduce the inductive bias of scanning strategies while ensuring that the model always maintains a good global receptive field.
+We adapt a simple approach of adding the spectrogram with the original feature map, allowing the model to scan simultaneously in both the frequency and spatial domains, thereby establishing a unified representation of visual information across both domains.
 Furthermore, we employ two trainable parameters to dynamically adjust the information intensity in the frequency and spatial domains, aiming to reduce semantic gaps.
 We also introduce additional local correlations during patch embedding by using overlapping convolutions.
-Our Vim-F(H) is motivated by a simple idea and is easy to implement.
-Extensive experiments have demonstrated that it effectively improves the performance of Vim, achieving competitive results compared to advanced ViTs, CNNs, and ViMs.
-The effectiveness of our method for ViMs with hybrid encoders has not been fully studied.
-Appendix A Appendix
-A.1 Ablation study
-Effect of frequency domain scanning.
-We evaluated the scanning strategy of Vim-F, and the experimental details are shown in Table 5.
-As mentioned earlier, position embeddings are crucial for Vim, and removing them leads to a 1.7% performance drop for Vim-Ti.
-However, the global receptive field brought by frequency domain scanning reduces the model’s dependence on positional embeddings, resulting in a smaller performance drop for Vim-Ti-F.
-When both models use positional embeddings, Vim-Ti-F outperforms Vim-Ti by 0.4%.
-Effect of convolutional stem.
-As shown in Table 5, further replacing the traditional patch embedding in Vim-Ti with our proposed convolutional stem leads to a 0.9% improvement in performance.
-This shows that introducing additional local correlation into the patch embedding is effective.
-Table 5: Ablation study of frequency domain scanning and patch embedding in Vim-Ti-F(H) on ImageNet-1K.
-Model Top-1 (%)
-Vim-Ti (without Position Embedding) 73.0%
-Vim-Ti (base) 74.7%
-Vim-Ti-F 74.2%
-Vim-Ti-F + Position Embedding 75.1%
-Vim-Ti-F(H) 76.0%
+Vim-F(H) enhances the performance of Vim-F by introducing linear attention, which mitigates the sequence order sensitivity of the Mamba encoder.
+Extensive experiments demonstrate that our work effectively improves the performance of Vim, with Vim-F and Vim-F(H) achieving competitive results.

### v3

@@
-Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain
+
 Abstract
-In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences.
-Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive.
+In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as Mamba deep learning models, have made significant progress in modeling long sequences.
+Compared to Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), Vision Mamba (ViM) methods have not yet achieved fully competitive performance.
 To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective.
@@
 We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains.
-Moreover, we question the necessity of position embedding in ViM and remove it accordingly in Vim-F, which helps to fully utilize the efficient long-sequence modeling capability of ViM.
-Finally, we redesign a patch embedding for Vim-F, leveraging a convolutional stem to capture more local correlations, further improving the performance of Vim-F.
-Code is available at: https://github.com/yws-wxs/Vim-F.
-Index Terms:
-Vision Mamba, State Space Model, Fast Fourier Transform, Scanning strategy
-I Introduction
+Moreover, considering that Mamba remains essentially a Recurrent Neural Network (RNN), we question the necessity of position embedding in ViM and remove it accordingly in Vim-F.
+Vim-F has good scalability.
+As far as we know, its variant Vim-F(CF) is the first ViM model to use a convolution-free ViM encoder.
+Another variant, Vim-F(H), introduces a linear attention mechanism.
+This reduces the model’s sensitivity to the input sequence and achieves better performance.
+1 Introduction
 Visual representation learning stands as a pivotal research topic within the realm of computer vision.
@@
 Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks.
-Except for the initial ”patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible.
+Except for the initial “patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible.
 Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism.
 However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
-Figure 1: The overview of the Vim model.
 Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training.
@@
 Therefore, Vim lacks the rationale for using positional embeddings, and thus our Vim-F does not employ positional embeddings.
-Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining overall linear complexity.
+Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining approximately linear complexity.
 We refer to this hybrid architecture as Vim-F(H).
-We outline our contributions below:
-  • Based on Vim [1], we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships.
-To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
-  • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it.
-Additionally, we question the necessity of position embedding in Vim, thus Vim-F does not utilize position embedding.
-  • The experimental results demonstrate that the simple Vim variant, Vim-F, achieves better performance while maintaining similar computational costs and parameters.
-Additionally, introducing linear attention to reduce Vim-F’s sensitivity to the order of input sequences can further enhance the model’s performance.
-II Related Works
-II-A State Space Models
+Finally, we explore the significance and feasibility of the convolution-free visual Mamba encoder and propose a proof-of-concept network, Vim-F(CF).
+Through extensive experiments on image classification, object detection, instance segmentation, and semantic segmentation tasks, the simple Vim variants, Vim-F(H) and Vim-F(CF), achieve better performance.
+2 Related Works
+State Space Models.
 In mathematics, State Space Models (SSMs) are usually expressed as linear ordinary differential equations (ODEs).
@@
 where $\bm{A}\in\mathbb{R}^{N\times N}$, $\bm{B}\in\mathbb{R}^{D\times N}$ and $\bm{C}\in\mathbb{R}^{N\times D}$.
-II-B Overview of Vim
+Discretization.
+The primary objective of discretization is to convert ODEs into discrete functions, allowing the model to align with the sampling frequency of the input signal for more efficient computation.
+Following the work [3], the continuous parameters ($\bm{A}$, $\bm{B}$) can be discretized by zero-order hold rule with a given sample timescale $\Delta\in\mathbb{R}^{N}$:
+$\bar{\bm{A}}=e^{\Delta\bm{A}},$ (2)
+$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(e^{\Delta\bm{A}}-\bm{I})\cdot\Delta\bm{B},$
+$\bar{\bm{C}}=\bm{C},$
+$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$
+$y(t)=\bar{\bm{C}}h(t),$
+where $\bm{\bar{A}}\in\mathbb{R}^{N\times N}$, $\bm{\bar{B}}\in\mathbb{R}^{D\times N}$ and $\bm{\bar{C}}\in\mathbb{R}^{N\times D}$.
+To improve computational efficiency, the iterative process described in equation 2 can be accelerated through parallel computing using global convolutional operations:
+$\begin{split}\bm{y}&=\bm{x}\circledast\bm{\overline{K}}\\ \text{with}\quad\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{L-1}\bm{\overline{B}}),\end{split}$ (3)
+[…]
+Selective State Space Models (S6).
+Mamba [2] improves the performance of SSM by introducing Selective State Space Models (S6), allowing the continuous parameters to vary with the input enhances selective information processing across sequences, which extend the discretization process by selection mechanism:
+$\bar{\bm{B}}$ $=s_{\bm{B}}(x),$ (4)
+$\bar{\bm{C}}$ $=s_{\bm{C}}(x),$
+$\Delta$ $=softplus\big(\text{Parameter}+s_{\bm{A}}(x)\big),$
+where $s_{\bm{B}}(x)$ and $s_{\bm{C}}(x)$ are linear functions that project input $x$ into an $N$-dimensional space, while $s_{\bm{A}}(x)$ broadens a $D$-dimensional linear projection to the necessary dimensions.
+Parameter represents a trainable parameter matrix.
+softplussoftplus is a mathematical operation defined as: $softplus(x)=log(1+e^{x})$.
+In this paper, we use the term state space model (SSM) to refer to variants of SSMs, including S6.
+Figure 1: The overview of the Vim model.
+Overview of Vim.
 As mentioned earlier, the effectiveness of our work is demonstrated through the improvement of Vim.
@@
 where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of blocks, and $\mathbf{Norm}$ is the normalization layer.
-Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains.
-Due to the translation invariance of the Fourier transform, the Fourier transforms of figures (a) and (b) are identical.
-III Rethinking the Scanning Strategy of ViMs
+The main difference between the Vim block and the classic Mamba block [2] resides in that Vim introduces a bidirectional scanning strategy, employing two independent branches to process the forward and backward sequences respectively.
+Each branch comprises a convolutional layer and an S6 module.
+3 Rethinking the Scanning Strategy of ViMs
 The non-causal nature of two-dimensional spatial patterns in images is inherently inconsistent with the causal processing framework of SSM.
 Therefore, it is meaningful to explore methods for serializing images.
-However, the majority of existing work on scanning strategies is typically based on a rather idealized assumption that introducing more or more complex inductive biases regarding the order of one-dimensional sequences can potentially enhance Mamba’s ability to understand local image correlations.
+However, most existing works on scanning strategies are typically based on a rather idealized assumption, that is, introducing more or more complex one-dimensional sequence scanning strategies may help enhance Mamba’s ability to understand local two-dimensional correlations in images.
+However, most existing works on scanning strategies are usually based on a rather idealized assumption.
+Scanning the same sequence multiple times in different orders can expand Mamba’s receptive field in different directions, thereby enhancing Mamba’s ability to understand local 2D correlations in images.
 Theoretically, regardless of the order, reducing the distance between adjacent tokens in one direction in two-dimensional space inevitably increases the distance between adjacent tokens in another direction, thus these strategies are essentially equivalent.
-Additionally, in the absence of global information serving as a ”frame of reference,” these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences.
-In practice, works involving scanning strategy research, including Vim [1], VMamba [2], LocalMamba [3], EfficientVMamba [4], and PlainMamba [9], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies.
+Additionally, in the absence of global information serving as a reference, these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences.
+In practice, most works involving scanning strategy research, including Vim [22], LocalMamba [5], EfficientVMamba [14], and PlainMamba [20], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies.
+VMamba [9] conducted experiments on different scanning strategies using ImageNet-1K, and the results indicate that most strategies show no significant performance differences.
 Zhu et al.
-[10] conduct a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images.
+[23] conducted a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images.
 Through extensive experiments on the LoveDA, ISPRS Potsdam, and ISPRS Vaihingen datasets, they demonstrate that no single scanning strategy outperforms others, regardless of their complexity or the number of scanning directions involved.
 In summary, an effective scanning strategy needs to maintain a global receptive field throughout the scanning process and demonstrate its effectiveness through analysis and comparative experiments.
-IV Methodology
-In this section, we first elaborate on the significance and benefits of frequency-domain scanning.
+4 Methodology
+In this section, we first elaborate on the advantages of frequency-domain scanning and the scanning strategy we proposed.
 Subsequently, we present the specific details of patch embedding in Vim-F.
-Finally, we discuss the composition of the Vim-F and the Vim-F(H) block.
-IV-A 2D Scanning in the Frequency Domain
-The flattening method used in Vim [1] disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens.
-We address this limitation by introducing frequency-domain scanning.
+We then discuss the composition of the Vim-F and Vim-F(H) blocks.
+Finally, we explore the significance and feasibility of the convolution-free visual Mamba encoder and propose a proof-of-concept network, Vim-F(CF).
+4.1 Scanning in the Frequency Domain
+As analyzed in Section 3, maintaining a global receptive field during the scanning process is crucial for enhancing Mamba’s ability to understand local 2D correlations in images through improved scanning strategies.
+We believe that introducing frequency-domain scanning is a simple and straightforward solution.
 Specifically, we first perform a Fourier transform on the feature map to obtain:
-$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}f(x,y)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)},$ (4)
+$F(u,v)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}f(x,y)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)},$ (7)
 where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively.
-$f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain.
-$|F(u,v)|\in\mathbb{R}^{H\times W}$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum.
-Fourier transform has translation invariance.
-In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function.
-This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Fig.
-2.
-The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy.
-Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy.
-IV-B Patch Embedding
+$f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain.
+From Eq.
+(7), it can be seen that each frequency-domain coefficient $F(u,v)$ is related to the values of all pixels in the feature map.
+$|F(u,v)|$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum.
+Obviously, scanning on the amplitude spectrum can ensure a global receptive field throughout the scanning process.
+In addition, the Fourier transform has translation invariance.
+Specifically, if a function is translated in the spatial domain, its amplitude spectrum is the same as that of the original function.
+A brief proof is provided below: According to Eq.
+(7), assuming the translated feature map is f(xa,yb)f(x-a,y-b), where $a$ and $b$ are the translation amounts.
+The Fourier transform of the translated feature map is:
+$F^{\prime}(u,v)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}f(x-a,y-b)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)}.$ (8)
+Let $x^{\prime}=x-a$, $y^{\prime}=y-b$, then $x=x^{\prime}+a$, $y=y^{\prime}+b$. Substituting into Eq. (8), we get:
+$F^{\prime}(u,v)$ $=\sum_{x^{\prime}}\sum_{y^{\prime}}f(x^{\prime},y^{\prime})e^{-j2\pi\left(\frac{u(x^{\prime}+a)}{H}+\frac{v(y^{\prime}+b)}{W}\right)}$ (9)
+$=e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}\sum_{x^{\prime}}\sum_{y^{\prime}}f(x^{\prime},y^{\prime})e^{-j2\pi\left(\frac{ux^{\prime}}{H}+\frac{vy^{\prime}}{W}\right)}$
+$=e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}F(u,v).$
+We know that the modulus of the phase factor $e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}$ is 1, so the amplitude spectrum of $F^{\prime}(u,v)$ is:
+$|F^{\prime}(u,v)|=|e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}F(u,v)|=|F(u,v)|.$ (10)
+This property makes the amplitude spectrum an ideal “reference object” during scanning, so simultaneous scanning in both the frequency and spatial domains is better than scanning in either domain alone.
+As shown in Fig.
+2, adding the amplitude spectrum to the original feature map introduces global information while reducing inductive biases from one-dimensional serialization, thereby enhancing Mamba’s ability to understand semantic objects.
+Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains.
+Due to the translation invariance of the Fourier transform, the amplitude spectra of figures (a) and (b) are identical.
+4.2 Patch Embedding
 In Vim and ViTs, the patching strategy typically involves using large convolutional kernels (such as kernel sizes of 14 or 16) with non-overlapping convolutions.
@@
 Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer.
-The details of the adaptation with the two base models [1] (Vim-Ti and Vim-S) are shown in Table I.
-TABLE I: Convolutional stem for patch embedding.
+The details of the adaptation with the two base models [22] (Vim-Ti and Vim-S) are shown in Table 1.
+Table 1: Convolutional stem for patch embedding.
 Output size
 Vim-Ti Vim-S
-Block1 56$\times$56 7$\times$7, 48, stride 4 7$\times$7, 48, stride 4
+Block1
+$\times56$
+$\times7$, 48, stride 4 $\times7$, 48, stride 4
 padding 3 padding 3
-Block2 28$\times$28 $\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
-Block3 14$\times$14 $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
+Block2
+$\times28$
+$\begin{bmatrix}\text{2$\times$2, 96, stride 2}\\ \text{1$\times$1, 96}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$
+Block3
+$\times14$
+$\begin{bmatrix}\text{2$\times$2, 192, stride 2}\\ \text{1$\times$1, 192}\end{bmatrix}$ $\begin{bmatrix}\text{2$\times$2, 384, stride 2}\\ \text{1$\times$1, 384}\end{bmatrix}$
 The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16.
@@
 Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035.
-IV-C Vim-F Block
-As mentioned earlier, the Vim-F block is an improvement over the Vim block, as shown in Fig.
-3.
+4.3 Vim-F Block
+As mentioned earlier, the Vim-F block is an improvement over the Vim [22] block, as shown in Fig.
+3(a).
 Specifically, a 2D FFT is performed on the input to obtain the spectrogram.
@@
 After the linear layer mixes the hidden dimensions of the feature map, the cross-domain information is sent to the Vim block for processing.
-Figure 3: Vim-F block
 Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation.
@@
 For a 2D-DFT with dimensions $(L,D)$, its complexity is $\Omega(L\cdot D\cdot logL\cdot logD)$.
-The feedforward network is implemented by a linear layer, with a complexity of $\Omega(L\cdot D^{2})$.
+The complexity of the linear layer is $\Omega(L\cdot D^{2})$.
 Therefore, compared to the Vim block, the additional overhead of the Vim-F block is approximately linearly related to the sequence length $L$.
-IV-D Vim-F(H) Block
+Considering that Vim [22] flattens the input into a sequence of length 196, the hidden state dimensions $D$ of the two variants Vim-Ti and Vim-S are 192 and 384 respectively, with a state expansion ratio of 2.
+These additional computational overheads are even significantly smaller than any single linear projection in Eq.
+(4).
+Figure 3: The architecture of the Vim-F block and the Vim-F(H) block.
+4.4 Vim-F(H) Block
 Observing Eq.
@@
 We integrate the Vim-F block with the linear attention [6] to maintain approximately linear complexity, which we term Vim-F(H), and its structure is depicted in Fig.
+3(b).
+Figure 4: Illustration of (a) Overall architecture of Vim-F(CF), (b) Vim block structure for reference, and (c) Proposed Vim-F(CF) block structure.
+4.5 Vim-F(CF): Is the convolution necessary in the ViM encoder?
+We know that Mamba [2] was originally designed for NLP, with its core module S6 being good at capturing long-range dependencies.
+Convolutions were introduced to compensate for the model’s limitations in short-range context modeling.
+However, the sequence lengths processed by ViMs are usually short.
+For example, in Vim [22] and PlainMamba [20], images are flattened into sequences with a length of 196.
+Thus, from this perspective, retaining convolutions in the ViM encoder is unnecessary.
+Due to the influence of traditional visual model design concepts and experimental results, the majority of existing ViM encoders retain convolutional layers.
+VMamba [9] points out that depthwise convolutions play an important role in forming a global effective receptive field (ERF).
+We believe that removing the convolutional layer and adopting frequency domain scanning can make the modeling approach of ViM closer to that of Transformer, thereby enhancing the scalability and transferability of ViM.
+We use convolution-free Vim-F blocks to build the proof-of-concept network Vim-F(CF), as shown in Fig.
 4.
-Figure 4: Vim-F(H) block
-TABLE II: Comparison of different backbones on ImageNet-1K classification.
+Vim-F(CF) follows a classic 4-stage structure to enable SSM to model visual features across different scales.
+Unlike common CNNs, Transformers, ViMs, and their hybrid models, Vim-F(CF) does not incorporate attention mechanisms or feed forward network (FFN) modules.
+5 Experiments
+This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection, instance segmentation, and semantic segmentation.
+The specific architectures of Vim-F and Vim-F(H) are as follows: We first replace the patch embedding module, and then only the first 6 Vim blocks (25% of the total number of blocks) are substituted with either Vim-F or Vim-F(H) blocks to balance performance and efficiency.
+The number of linear attention heads in Vim-F(H) is set to 24.
+Table 2: Comparison of different backbones on ImageNet-1K classification.
 Method Params (M) FLOPs (G) Top-1 (%)
@@
 LocalVim-T [5] 8 1.5 76.5
-Vim-Ti-F (ours) 8 1.8 76.7
+Vim-Ti-F (ours) 7 1.5 76.7
 PlainMamba-L1 [20] 7 3.0 77.9
@@
 Vim-S [22] 26 5.1 80.5
-Vim-S-F (ours) 28 6.1 80.9
+Vim-F(CF) (ours) 22 5.2 80.7
+Vim-S-F (ours) 28 5.5 80.9
 LocalVim-S[5] 28 4.8 81.0
 Swin-T [10] 29 4.5 81.3
+PlainMamba-L2 [20] 25 8.1 81.6
 Vim-S-F(H) (ours) 37 6.8 81.6
 EfficientVMamba-B [14] 33 4.0 81.8
-VMamba-T [2] 22 5.6 82.2
-V Experiments
-This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation.
-The specific architectures of Vim-F and Vim-F(H) are as follows: We replaced only the first 6 Vim blocks (25% of the total number of blocks) with either Vim-F or Vim-F(H) blocks to balance performance and efficiency while maintaining approximately the same number of parameters.
-The number of linear attention heads in Vim-F(H) is set to 24.
-V-A Image Classification
-Settings.
+VMamba-T [9] 30 4.9 82.6
+5.1 Image Classification
+5.1.1 Settings.
 We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set.
@@
 Therefore, we adjust the total batch size and the initial learning rate to 384 and $3.75\times 10^{-4}$ respectively according to the linear scaling rule[1].
-Results.
-We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table II.
+5.1.2 Results.
+We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table 2.
 Vim-F outperforms the original Vim in terms of performance, and Vim-F(H) further enhances Vim-F’s performance by introducing linear attention.
-It is worth noting that SOTA ViMs such as VMamba [2], PlainMamba [9], and EfficientVMamba [4] all adopt more complex CNN-Mamba hybrid encoding architectures, and their network structures are similar to advanced convolutional networks like ConvNeXt [19].
-Since these works have not conducted the necessary comparative experiments to address this issue, it is reasonable to suspect that the excellent performance of these ViMs may primarily stem from the use of depthwise separable convolutions.
-However, Vim-F still uses a pure SSM encoder to model visual information.
+Although not meticulously designed, the proof-of-concept network Vim-F(CF) matches Vim-S in performance and computational complexity and has a parameter count advantage, indicating that convolution-free ViM encoders and associated model architectures are also a promising research direction.
+It is worth noting that SOTA ViMs such as VMamba [9], PlainMamba [20], and EfficientVMamba [14] all adopt more complex CNN-Mamba hybrid encoders, using depthwise separable convolutions to extract features in a 2D space.
+In addition, their network architectures also draw on advanced networks like ConvNeXt [11], MetaFormer [21], MobileViT [13].
+Since these works have not conducted ablation experiments on the necessity of the SSM block, it is reasonable to suspect that the exceptional performance of these ViMs may mainly be attributed to the introduction of depthwise separable convolutions and advanced network architectures.
+Unlike these works, Vim-F and Vim-F(H) do not use depthwise separable convolutions and largely follow the design philosophy of Vim’s network architecture.
 LocalVim[5] is one of the representative works that improve the scanning strategy for Vim.
@@
 In comparison, our works achieve a better balance between performance, efficiency, and model complexity.
-TABLE III: Object detection and instance segmentation results on COCO.
+Table 3: Ablation study on the proposed patch embedding module and Vim-F block proportion.
+Method Params (M) FLOPs (G) Top-1 (%)
+Vim-Ti [22] 7 1.5 76.1
+- Position Embedding 7 1.5 74.8
+Vim-Ti-F-V0 7 1.5 76.3
+Vim-Ti-F-V1 7 1.5 76.7
+Vim-Ti-F-V2 8 1.7 76.8
+Vim-Ti-F-V3 9 1.9 77.0
+5.1.3 Ablation Study on ImageNet-1K.
+Here, taking Vim-Ti-F as an example, we conduct an ablation study on the effectiveness of our proposed patch embedding module and the impact of the Vim-F block proportion.
+We configured four variants of Vim-Ti-F, namely V0 to V3, with Vim-F block proportions set at 0%, 25%, 50%, and 100%, respectively.
+As shown in Table 3, removing position embedding caused a 1.3% performance degradation of Vim-Ti.
+Vim-Ti-F-V0 showed slightly better performance than Vim-Ti, indicating that using a convolutional stem with overlapping convolutions as patch embedding is a more effective approach.
+Additionally, substituting more Vim blocks with Vim-F blocks can lead to steady performance improvement.
+However, due to the overall representation limitations of the model, it is unnecessary to use Vim-F blocks exclusively.
+Table 4: Object detection and instance segmentation results on COCO.
+Backbone Params FLOPs APb $^{\mathrm{b}}_{50} AP$$^{\mathrm{b}}_{75}$ APm $^{\mathrm{m}}_{50} AP$$^{\mathrm{m}}_{75}$
+ResNet-18 [4] 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
+Vim-Ti [22] 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
+Vim-Ti-F (ours) 28M 189G 37.7 60.8 40.9 35.7 57.8 38.0
+ResNet-50 [4] 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
+Vim-Ti-F(H) (ours) 28M 189G 39.3 61.8 42.6 36.7 58.9 39.2
+EfficientVMamba-S [14] 31M 197G 39.5 60.8 43.1 35.8 57.4 38.3
+ResNet-101 [4] 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
+Vim-S [22] 44M 272G 40.9 63.9 45.1 37.5 59.4 40.2
+Vim-S-F (ours) 47M 273G 41.6 62.7 45.6 37.9 60.8 40.7
+Vim-F(CF) (ours) 42M 272G 42.0 62.5 45.9 38.4 60.9 41.3
+Swin-T [10] 48M 267G 42.7 65.2 46.8 39.2 62.1 42.1
+Vim-S-F(H) (ours) 57M 275G 42.9 65.4 47.1 39.3 62.2 42.2
+5.2 Object Detection and Instance Segmentation
+5.2.1 Settings.
 […]
-ResNet-18 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
-Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
-PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
-Vim-F-Ti (ours) 28M 189G 37.7 60.8 40.9 35.7 57.8 38.0
-ResNet-50 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
-Vim-F(H)-Ti (ours) 28M 189G 38.2 61.8 41.7 35.9 58.6 38.0
-EfficientVMamba-S 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
-ResNet-101 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
-Vim-S 44M 272G 40.9 63.9 45.1 37.9 60.8 40.7
-Vim-F-S (ours) 47M 273G 41.6 64.1 45.5 38.4 60.9 41.3
-Swin-T 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
-Vim-F(H)-S (ours) 57M 275G 42.9 65.4 47.1 39.2 62.1 42.1
-Swin-S 69M 354G 44.8 66.6 48.9 40.9 63.2 44.2
-V-B Object Detection and Instance Segmentation
+5.2.2 Results.
+We summarize the comparison results of Vim-F, Vim-F(H), and Vim-F(CF) with other backbones in Table 4.
+It can be seen that our Vim-F consistently outperforms Vim.
+Notably, Vim-S-F outperforms Vim-F(CF) in classification tasks, while Vim-F(CF) surpasses Vim-S-F in downstream dense prediction tasks.
+Similar to the results in classification tasks, Vim-F(H) exhibits better performance compared to both Vim-F and Vim-F(CF), and it is competitive with state-of-the-art ViT architectures.
+5.3 Semantic Segmentation
 Settings.
-We use Mask-RCNN as the detector to evaluate the performance of the proposed Vim-F for object detection and instance segmentation on the MSCOCO 2017 dataset.
-Since the backbone of Vim-F is non-hierarchical, following ViTDet[20], we only used the last feature map from the backbone and generated multi-scale feature maps through a set of convolutions or deconvolutions to adapt to the detector.
-The remaining settings were consistent with Swin[17].
-Specifically, we employ the AdamW optimizer and fine-tune the pre-trained classification models (on ImageNet-1K) for both 12 epochs (1$\times$ schedule).
-The learning rate is initialized at $1\times 10^{-4}$ and is reduced by a factor of 10×10$\times$10 × at the 9th and 11th epochs.
+Following Vim [22], we train UperNet [18] with our models on ADE20K dataset.
+In training, we employ AdamW with a weight decay of 0.01, and a total batch size of 16 to optimize models.
+The employed training schedule uses an initial learning rate of $6\times 10^{-5}$, linear learning rate decay, a linear warmup of 1500 iterations, and a total training of 160K iterations.
+Table 5: Results of semantic segmentation on ADE20K.
+Backbone Params mIoU
+Vim-Ti [22] 34M 40.1
+Vim-Ti-F (ours) 34M 40.4
+ResNet-50 [4] 67M 40.7
+Vim-S [22] 57M 43.9
+Vim-S-F (ours) 58M 44.2
+Vim-F(CF) (ours) 55M 44.4
+Swin-T [10] 60M 44.5
+Vim-S-F(H) (ours) 47M 44.8
 Results.
-We summarize the comparison results of Vim-F and Vim-F(H) with other backbones in Table III.
-It can be seen that our Vim-F consistently outperforms Vim.
-Similar to the results in classification tasks, Vim-F(H) exhibits better performance than Vim-F and is competitive with advanced ViT architectures.
-VI Conclusion
+The results are presented in Table 5.
+Similar to the experimental results on MSCOCO, Vim-F and its variants demonstrate superior performance to Vim [22], ResNet [4], and Swin[10] in the semantic segmentation task, which further confirms the effectiveness of our work.
+6 Conclusion
 One of the main challenges currently faced by ViMs is the destruction of local two-dimensional dependency relationships when spatial data is flattened into one-dimensional tokens due to the non-causal nature of visual data.
@@
 Vim-F(H) enhances the performance of Vim-F by introducing linear attention, which mitigates the sequence order sensitivity of the Mamba encoder.
-Extensive experiments demonstrate that our work effectively improves the performance of Vim, with Vim-F and Vim-F(H) achieving competitive results.
+We constructed a proof-of-concept network, Vim-F(CF), and found that even after removing the convolutional layers from Vim-F, the model can still effectively model visual information.
+Extensive experiments show that our work effectively boosts Vim’s performance.
+Vim-F, Vim-F(H), and Vim-F(CF) achieve results competitive with advanced ViM models.